OpenStudy (anonymous):

Write 806,000,000 in scientific notation.

OpenStudy (anonymous):

8.06*10^8

OpenStudy (ash2326):

We have \[806,000,000\] Let's count the zeros=> we have 6 zeros here. So intermediate form is \[806 \times 10^6\] now 806= 8.06 * 100 so \[8.06 \times 100 \times 10^6\] So the scientific form is \[8.06*10^8\]

OpenStudy (anonymous):

Thanks btw , how do you type exponents on the keyboard ?

OpenStudy (anonymous):

806,000,000 in scientific notation is 8.06 x 10^8. & @ZachKetchup Hold shift and press 6